# N-GLARE: An Non-Generative Latent Representation-Efficient LLM Safety Evaluator

### v1

@@ Abstract
Evaluating the safety robustness of LLMs is critical for their deployment. However, mainstream Red Teaming methods rely on online generation and black-box output analysis. These approaches are not only costly but also suffer from feedback latency, making them unsuitable for agile diagnostics after training a new model. To address this, we propose N-GLARE (A Non-Generative, Latent Representation-Efficient LLM Safety Evaluator). N-GLARE operates entirely on the model’s latent representations, bypassing the need for full text generation. It characterizes hidden layer dynamics by analyzing the APT (Angular-Probabilistic Trajectory) of latent representations and introducing the JSS (Jensen-Shannon Separability) metric. Experiments on over 40 models and 20 red teaming strategies demonstrate that the JSS metric exhibits high consistency with the safety rankings derived from Red Teaming. N-GLARE reproduces the discriminative trends of large-scale red-teaming tests at less than 1% of the token cost and the runtime cost, providing an efficient output-free evaluation proxy for real-time diagnostics.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have developed rapidly, with their capability boundaries continuously expanding. They have been deployed in a wide range of application scenarios, exerting substantial impact on social production and daily life [grattafiori2024llama, zhou2024large]. However, alongside this growth in capability, security risks and alignment issues have become increasingly prominent, emerging as a critical bottleneck for large-scale, safety-critical deployment [das2025security, dong2024attacks]. Trained on vast and often uncurated Internet corpora, LLMs can inadvertently internalize social biases and potentially harmful knowledge. Without robust safety alignment, they may produce unsafe outputs when exposed to malicious prompts, complex contexts, or specific triggers—for example, generating cyberattack code [wei2023jailbroken], spreading misleading information [russinovich2025great], or guiding embodied agents toward physically harmful behaviors [yang2025concept]. Ensuring that LLMs remain aligned with human values and safety norms is therefore a central challenge.
@@
Evaluating model safety is a foundational step toward achieving such alignment. The dominant paradigm in both industry and academia is currently Red Teaming [lin2025against], which simulates adversarial user behavior by designing diverse, subtle attack prompts to probe the model and elicit unsafe outputs. While this online, behavior-based testing has proved effective for discovering vulnerabilities, it suffers from two fundamental limitations. First, it is highly resource-intensive: test coverage depends on the scale and diversity of attack prompts, which require continual human and computational investment as new attack vectors emerge. Second, Red Teaming is essentially a black-box methodology. It only observes input–output behavior and provides little insight into the model’s internal safety state. A refusal that looks safe at the surface level may simply be triggered by brittle pattern matching, rather than reflecting a robust internal representation of harmfulness.
@@
To move beyond purely behavioral stress testing, recent work has begun to analyze LLMs in their latent representation space [zou2023representation, li2025revisiting]. Representation engineering shows that when a model processes different concepts (including abstract notions such as “truthfulness,” “refusal,” or “harmfulness”), its hidden activations form structured, often linearly separable patterns. Let $h_{l,p}\in\mathbb{R}^{d}$ denote the activation at layer $l$ and position $p$. For two contrasting semantic sets $\mathcal{S}^{+}$ and $\mathcal{S}^{-}$, their mean activations differ along concept directions of the form
$o\approx\mathbb{E}[H(\mathcal{S}^{+})]-\mathbb{E}[H(\mathcal{S}^{-})],$ (1)
@@
suggesting that high-level semantics are explicitly encoded in latent geometry. Existing work primarily uses such directions at the prompt level, for example to steer individual generations or to classify the risk of a specific input. However, these techniques are typically model-specific and do not directly provide a model-level safety indicator that is comparable across architectures and checkpoints.
@@
In this paper, we take a step in that direction and ask: can we evaluate model safety without generating full outputs, solely by examining how its internal representations evolve under different safety conditions? To this end, we propose N-GLARE (A Non-Generative, Latent Representation-Efficient LLM Safety Evaluator), a framework that operates entirely in latent space and bypasses token-level generation. For each model, we construct Angular-Probabilistic Trajectories (APT) of hidden representations under different probing conditions: benign ($B$), jailbreak ($J$), plainquery ($P$) and an idealized refusal counterfactual ($R$). After normalizing these trajectories along a standardized progress axis, we model benign activations as a low-dimensional manifold and define a geometric turning angle that measures how strongly a trajectory deviates away from this benign manifold over time. We then introduce Jensen–Shannon Separability (JSS), which compares the slice-wise distributions of these turning angles across conditions, and aggregates them into model-level safety indicators, introducing other JSS-based safety scores.
Figure 1: JSS Shows Stronger Safety Alignment in Qwen3-4B-SafeRL Compared to Base and Alignment-Removed Variants. Statistics on the JSS distribution for 100 jailbreak trajectories and the ideal refusal pattern.
@@
Figure 1 illustrates the core intuition: when comparing multiple variants of the same base model (e.g., RL-aligned, base, and alignment-removed versions of Qwen3-4B), we observe that better-aligned variants exhibit more pronounced geometric separation between benign and jailbreak trajectories in latent space, and their jailbreak trajectories are also further away from the ideal refusal pattern. This suggests that internal trajectory separability can serve as a proxy for the underlying safety state of the model, potentially reproducing the discriminative power of large-scale Red Teaming at a fraction of the cost.
@@
We formalize this idea in Section 3 by casting safety evaluation as a problem of comparing representation trajectories under $(B,J,R,P)$ conditions. We then instantiate our method in Section 4: benign activations define a whitened manifold, the trajectory turning angle encodes instantaneous deviation from this manifold, and Jensen–Shannon divergence over angle distributions along the progress axis yields JSS metrics that are both geometrically interpretable and amenable to cross-model comparison. Experiments on over 40 models and 20 red-teaming strategies demonstrate that N-GLARE faithfully reproduces the safety rankings induced by traditional Red Teaming while using less than 1% of the token and runtime cost, making it suitable for agile diagnostics after training new models or updating safety mechanisms.
@@ 2.1 LLM Security and Red Teaming
Despite the widespread implementation of safety alignment in contemporary mainstream LLMs, they continue to face extensive security challenges in practical application [Cui2024Risk]. Attackers frequently employ meticulously designed adversarial inputs to bypass the model’s safety guardrails [zhu2023promptbench, paulus2024advprompter]. Through so-called “jailbreak attacks,” models are induced to circumvent their established policies, consequently generating harmful content. Such content may be exploited for academic misconduct [oliveira2025human], facilitating cyberattacks [xu2024autoattacker], or infringing upon copyright [xu2024llms].
Consequently, ensuring that model outputs adhere to ethical norms and effectively mitigating potential risks have become indispensable research directions in the field of artificial intelligence. In this context, the security evaluation of large models is paramount, wherein Red Teaming has emerged as a core methodology for assessing model safety.
Red Teaming is an adversarial evaluation method centered on proactively discovering and identifying potential model vulnerabilities by simulating attacker strategies and behaviors. Existing Red Teaming approaches can be broadly categorized into two types. The first focuses on content safety in single-turn interactions, such as leveraging algorithms like Monte Carlo Tree Search (MCTS) and evolutionary algorithms to optimize attack prompts [sadasivan2024fast], or generating jailbreak templates through the mutation of seed prompts [yu2023gptfuzzer]. The second category addresses security risks in multi-turn dialogue scenarios, typically employing specialized agents for attack planning, executing dynamic multi-turn jailbreaks, and evaluating the attack’s efficacy [rahman2025x, weng2025foot, ren2025llms].
Although existing methods show a trend towards diversification, their mainstream paradigm still generally follows a three-stage process: first, constructing the attack prompt; second, obtaining the target model’s output for the prompt; and finally, judging the success of the attack based on the evaluation of the model’s output results. This model leads to particularly high evaluation overhead for Red Teaming.
@@ 2.2.1 Embedding Analysis
Embedding analysis reveals semantic meaning by examining the static geometry of latent space vectors. Common methods like cosine similarity ($Sim(\vec{A},\vec{B})=\frac{\vec{A}\cdot\vec{B}}{\|\vec{A}\|\|\vec{B}\|}$) and dimensionality reduction visualization like t-SNE can reveal semantic clusters (e.g., “king” with “queen”) [rogers2020primer]. Furthermore, domain-adaptation analysis shows that fine-tuning (e.g., BioBERT [lee2020biobert]) reshapes the latent space, forming more refined domain-specific subspaces to enhance downstream task performance.
@@ 2.2.2 Probing and Intervention
This technique shifts from passive analysis to active decoding and control. Early probing [tenney2019bert] used simple linear classifiers ($y=f(H)$) to test if latent activations $H$ encoded specific linguistic properties, discovering that information is stored hierarchically in LLMs.
A more advanced paradigm, Representation Engineering (RepE) [zou2023representation], controls model behavior by directly manipulating latent representations. Its core is identifying vector directions, or concept operators $o^{c}$, that represent high-level concepts (e.g., truthfulness). This is often achieved by contrasting the activation differences from different input sets ($\mathcal{S}^{c+}$ vs $\mathcal{S}^{c-}$):
$o^{c}\approx\mathbb{E}[H(\mathcal{S}^{c+})]-\mathbb{E}[H(\mathcal{S}^{c-})]$ (2)
Here, $H(\mathcal{S})$ represents the mean activation of model $\mathcal{M}$ when processing a specific input set $\mathcal{S}$. RepE is widely applied in LLM safety [cai2024self, wang2025model, wang2024inferaligner, cao2024nothing], for instance, by using positive or negative steering vectors to enhance truthfulness [marks2023geometry, li2023inference, wang2025adaptive, rimsky2024steering] or suppress meta-behaviors like “refusal” [arditi2024refusal].
2.2.3 Attention Mechanism Analysis
Attention mechanism analysis interprets the model’s information flow via its weight distribution $A=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)$. Here, the $Q,K,V$ matrices are linear projections of the layer input $h_{l-1}$. Heatmap visualization [vig2019multiscale] reveals that attention exhibits hierarchical properties (from syntax to semantics across layers). At the same time, research [michel2019sixteen] indicates high redundancy in attention mechanisms. To address its computational complexity, models like Perceiver [jaegle2021perceiver] introduce low-dimensional latent variables to compress the computation.
@@ 2.2.4 Latent Space Dynamic Analysis
Dynamic analysis focuses on the activation trajectories during the reasoning process. Researchers not only track attention changes during Chain-of-Thought (CoT) [wei2022chain] but also begin to analyze the latent space trajectories themselves. For example, Li et al. [li2025core] proposed Chain-of-Reasoning Embedding (CoRE), which treats the reasoning process as a series of hidden states $h_{l,p_{T}}$. They found that redundant reasoning (”overthinking”) manifests as identifiable cyclical fluctuations in the trajectory. Based on this, the CoRE-Eval framework was developed, enabling label-free self-evaluation to dynamically terminate inefficient reasoning, thereby improving efficiency.
@@
3 Problem Statement
Safety evaluation of large language models is currently dominated by generative red-teaming: a model is queried with adversarial prompts generated by red-teaming LLMs, asked to produce multi-token outputs, and then judged by a human or a separate LLM-based evaluator. This approach is expensive (because evaluation cost grows with generated tokens), noisy (because harmfulness must be inferred from surface text), and incomplete (because it reveals only the final behavior, not the underlying safety state).
Formally, let $M_{\text{subject}},M_{s}$ be a model and let $x$ be a potentially harmful input drawn from an adversarial distribution $\mathcal{D}_{\mathrm{adv}}$ or online $M_{\text{attacker}},M_{r}$. A generative evaluator, another LLM $M_{\text{evaluator}},M_{e}$ observes the decoded output $y=M_{s}(x)$ and produces an assessment $\mathrm{Eval_{M_{e}}}(y)$ that approximates the latent safety function $f_{\mathrm{safety}}(M_{s},x)$. That is,
$\mathrm{Eval_{M_{e}}}(y)\approx f_{\mathrm{safety}}(M_{s},x),$
where this 0/1 function, $f_{\mathrm{safety}}$ represents whether the model processes $x$ in a safe or unsafe manner. This evaluation pipeline is entirely output-dependent (by analyzing the natural language text) and therefore cannot access $f_{\mathrm{safety}}(M_{s},x)$ directly.
In contrast, a non-generative evaluator also seeks to estimate $f_{\mathrm{safety}}(M_{s},x)$ without decoding any output tokens. Instead, it relies only on information available inside the $M_{s}$. The goal is to design a mapping
$g:(M_{s},x)\;\mapsto\;f_{\text{safety}}(M_{s},x),$
where $g(M,x)$ is a model-level safety signal that: (i) approximates the same generative-based quantity $\mathrm{Eval_{M_{e}}}(y)$ evaluated by LLM-based evaluator, (ii) is comparable across models and prompts, and (iii) does not rely on online red-teaming text generation or attackers $M_{r}$.
The core problem addressed in this work is therefore:
How can we construct a non-generative safety metric $g(\cdot,\cdot)$ that reliably reflects a model’s internal safety state and correlates with its behavioral robustness observed under traditional red-teaming?
This paper develops such a metric by operating entirely in the representation space, enabling low-cost, model-level safety evaluation without generation.
@@ 4 Methodology
We now instantiate the representation-space safety evaluation framework outlined above. Given a set of models $\mathcal{M}=\{M_{1},\dots,M_{K}\}$ and probing conditions $A\in\{B,J,R,P\}$ (benign, jailbreak, ideal refusal), our goal is to map internal activation trajectories to scalar safety indicators. The pipeline proceeds in three stages: (i) extracting layer-wise hidden states and normalizing them along a standardized progress axis; (ii) fitting a benign manifold and defining a geometric turning angle that measures instantaneous deviation from this manifold; and (iii) aggregating slice-wise Jensen–Shannon divergences of turning-angle distributions into Jensen–Shannon Separability (JSS) metrics and a normalized safety score.
4.1 Trajectory Extraction and Standardized Progress
For each model $M$ and probing condition $A\in\{B,J,R,P\}$ (defined in Section 4.3), we construct probing sequences that form either genuine interaction histories (e.g., multi-turn dialogues) or canonical prefix expansions for single-turn prompts. At each node $t$ in such a sequence, we identify a fixed decision-making cross-section, chosen in our experiments as the point immediately before the model generates the next assistant response. A single forward pass at node $t$ yields the last-token hidden activations across all transformer layers,
@@ 4.1 Trajectory Extraction and Standardized Progress
where $L$ is the number of layers and $d_{\ell}$ is the hidden dimension at layer $\ell$.
To improve cross-model comparability and reduce dimensionality, we partition layers into a small number of layer groups $G\in\{\text{lower},\text{middle},\text{upper}\}$ and average within each group, obtaining group-level representations
$h_{t}^{(A,G)}\in\mathbb{R}^{d},$ (4)
with a common dimension $d$. For each condition $A$ and group $G$, the sequence $\{h_{t}^{(A,G)}\}_{t=1}^{T_{A}}$ defines a trajectory in latent space.

### v2

@@ Abstract
Evaluating the safety robustness of LLMs is critical for their deployment. However, mainstream Red Teaming methods rely on online generation and black-box output analysis. These approaches are not only costly but also suffer from feedback latency, making them unsuitable for agile diagnostics after training a new model. To address this, we propose N-GLARE (A Non-Generative, Latent Representation-Efficient LLM Safety Evaluator). N-GLARE operates entirely on the model’s latent representations, bypassing the need for full text generation. It characterizes hidden layer dynamics by analyzing the APT (Angular-Probabilistic Trajectory) of latent representations and introducing the JSS (Jensen-Shannon Separability) metric. Experiments on over 40 models and 20 red teaming strategies demonstrate that the JSS metric exhibits high consistency with Red Teaming safety rankings at less than 1% token and runtime cost.
N-GLARE: An Non-Generative Latent Representation-Efficient LLM Safety Evaluator
Zheyu Lin1, Jirui Yang2, Yukui Qiu3 Hengqi Guo2, Yubing Bao2, Yao Guan2 1University of California, Riverside, 2Fudan University, 3South China University of Technology Zheyu Lin: 1298lzy@gmail.com Jirui Yang: yangjr23@m.fudan.edu.cn Yukui Qiu: 202162311356@mail.scut.edu.cn Hengqi Guo: hqguo22@m.fudan.edu.cn Yubing Bao: ybbao23@m.fudan.edu.cn Yao Guan: yguan25@m.fudan.edu.cn
@@ 1 Introduction
Large Language Models must be aligned to ensure their behavior consistently follows human values and safety norms, rather than merely maximizing task performance Das et al. (2025); Dong et al. (2024). Without such alignment, models may generate harmful, biased, or unsafe outputs when facing adversarial prompts or complex contexts Wei et al. (2023); Russinovich et al. (2025); Yang et al. (2025).
@@
Evaluating model safety is a foundational step toward achieving such alignment. The dominant paradigm in both industry and academia is currently Red Teaming Lin et al. (2025), which simulates adversarial user behavior by designing diverse, subtle attack prompts to probe the model and elicit unsafe outputs. While this online, behavior-based testing has proved effective for discovering vulnerabilities, it suffers from two fundamental limitations. First, it is highly resource-intensive: test coverage depends on the scale and diversity of attack prompts, which require continual human and computational investment as new attack vectors emerge. Second, Red Teaming is essentially a black-box methodology. It only observes input–output behavior and provides little insight into the model’s internal safety state. A refusal that looks safe at the surface level may simply be triggered by brittle pattern matching, rather than reflecting a robust internal representation of harmfulness.
@@
To move beyond purely behavioral stress testing, recent work analyzes LLMs in their latent representation space Zou et al. (2023); Li et al. (2025b). Representation engineering shows that different semantic concepts—such as truthfulness, refusal, or harmfulness—correspond to structured patterns in hidden activations, and contrasting concept sets can be distinguished by consistent directions in this space, suggesting that high-level semantics are explicitly encoded in latent geometry.
@@
Existing representation analysis work primarily uses such directions at the prompt level, for example to steer individual generations or to classify the risk of a specific input. However, these techniques are typically model-specific and do not directly provide a model-level safety indicator that is comparable across architectures and checkpoints.
In this paper, we take a step in that direction and ask: Can we evaluate model safety without generation, using internal representations that correlate with red-teaming robustness?
@@
To this end, we propose N-GLARE (A Non-Generative, Latent Representation-Efficient LLM Safety Evaluator). For each model, we construct Angular-Probabilistic Trajectories (APT) of hidden representations under different probing conditions: benign ($B$), jailbreak ($J$), plainquery ($P$) and an idealized refusal counterfactual ($R$). After normalizing these trajectories along a standardized progress axis, we model benign activations as a low-dimensional manifold and define a geometric turning angle that measures how strongly a trajectory deviates away from this benign manifold over time. We then introduce Jensen–Shannon Separability (JSS), which compares the slice-wise distributions of these turning angles across conditions, and aggregates them into model-level safety indicators, introducing other JSS-based safety scores.
Figure 1: JSS exhibits clear separability across safety states while achieving orders-of-magnitude higher token efficiency than red-teaming baselines.
The core intuition is that: internal trajectory separability can serve as a proxy for the underlying safety state of the model, potentially reproducing the discriminative power of large-scale Red Teaming at a fraction of the cost.
@@
Figure 1 supports the core intuition: when comparing multiple variants of the same base model (e.g., RL-aligned, base, and safety-removed versions of Qwen3-4B), we observe that better-aligned variants exhibit more pronounced geometric separation between different trajectories in latent space.
@@
Experiments on over 40 models and 20 red-teaming strategies demonstrate that N-GLARE faithfully reproduces the safety rankings induced by traditional Red Teaming while using less than 1% of the token and runtime cost, making it suitable for agile diagnostics after training new models or updating safety mechanisms.
@@ 2.1 LLM Security and Red Teaming
Red teaming evaluates LLM safety by simulating attacker strategies to proactively identify and probe model vulnerabilities.Existing Red Teaming approaches can be broadly categorized into two types. The first focuses on content safety in single-turn interactions, such as leveraging algorithms like Monte Carlo Tree Search (MCTS) and evolutionary algorithms to optimize attack prompts Sadasivan et al. (2024), or generating jailbreak templates through the mutation of seed prompts Yu et al. (2023). The second category addresses security risks in multi-turn dialogue scenarios, typically employing specialized agents for attack planning, executing dynamic multi-turn jailbreaks, and evaluating the attack’s efficacy Rahman et al. (2025); Weng et al. (2025); Ren et al. (2025).
Existing methods, despite increasing diversification, largely follow a three-stage pipeline that includes attack prompt construction, target model generation, and output-based evaluation, which inherently incurs high overhead in red-teaming assessments.
@@ 2.2.1 Embedding Analysis
Embedding analysis reveals semantic meaning by examining the static geometry of latent space vectors. Common methods like cosine similarity ($Sim(\vec{A},\vec{B})=\frac{\vec{A}\cdot\vec{B}}{\|\vec{A}\|\|\vec{B}\|}$) and dimensionality reduction visualization like t-SNE can reveal semantic clusters (e.g., “king" with “queen") Rogers et al. (2020). Furthermore, domain-adaptation analysis shows that fine-tuning such as BioBERT Lee et al. (2020) reshapes the latent space, forming more refined domain-specific subspaces to enhance downstream task performance.
@@ 2.2.2 Probing and Intervention
This technique shifts from passive analysis to active decoding and control. Early probing Tenney et al. (2019) used simple linear classifiers ($y=f(H)$) to test if latent activations $H$ encoded specific linguistic properties, discovering that information is stored hierarchically in LLMs.
A more advanced paradigm, Representation Engineering (RepE) Zou et al. (2023), controls model behavior by directly manipulating latent representations. Its core is identifying vector directions, or concept operators $o^{c}$, that represent high-level concepts (e.g., truthfulness). This is often achieved by contrasting the activation differences from different input sets ($\mathcal{S}^{c+}$ vs $\mathcal{S}^{c-}$): $o^{c}\approx\mathbb{E}[H(\mathcal{S}^{c+})]-\mathbb{E}[H(\mathcal{S}^{c-})]$ Here, $H(\mathcal{S})$ represents the mean activation of model $\mathcal{M}$ when processing a specific input set $\mathcal{S}$. RepE is widely applied in LLM safety Cai et al. (2024); Wang et al. (2025a, 2024); Cao et al. (2024), for instance, by using positive or negative steering vectors to enhance truthfulness Marks and Tegmark (2023); Li et al. (2023); Wang et al. (2025b); Rimsky et al. (2024) or suppress meta-behaviors like “refusal" Arditi et al. (2024).
@@ 2.2.3 Latent Space Dynamic Analysis
Dynamic analysis focuses on the activation trajectories during the reasoning process. Researchers not only track attention changes during Chain-of-Thought (CoT) Wei et al. (2022) but also begin to analyze the latent space trajectories themselves. For example, Li et al. Li et al. (2025a) proposed Chain-of-Reasoning Embedding (CoRE). They found that redundant reasoning manifests as identifiable cyclical fluctuations in the trajectory. Based on this, it enables label-free self-evaluation to dynamically terminate inefficient reasoning, thereby improving efficiency.
@@ 3 Methodology
This section formalizes the representation-space safety evaluation framework proposed above. Our goal is to map internal activation trajectories to scalar safety metrics. We first define the notation and probing conditions, followed by the framework’s three core stages: (i) trajectory extraction and standardization; (ii) benign manifold fitting and geometric turning angle calculation; and (iii) distribution divergence measurement and safety score aggregation.
3.1 Preliminaries: Symbols and Probing Conditions
Given a model set $\mathcal{M}=\{M_{1},\dots,M_{K}\}$, we analyze internal representation behaviors under varying inputs. We define the probing condition $A\in\{B,J,R,P\}$ to correspond to four interaction history types:
Benign ($B$): Standard task instructions conforming to natural distributions, used to fit the benign reference manifold.
Jailbreak ($J$): Adversarial sequences designed to induce harmful outputs.
Ideal Refusal ($R$): Identical to $J$ in user content (retaining harmful intent) but forcing the assistant response to a policy-compliant refusal template. This represents the model’s ideal defensive state against attacks.
Plain Query ($P$): Direct presentation of harmful intent (without disguise) to capture the model’s unconstrained internal response to harmful information.
To enhance cross-model comparability and reduce dimensionality, we partition Transformer layers into groups $G\in\{\text{lower},\text{middle},\text{upper}\}$ and average within each group to obtain group-level representations $h^{(A,G)}$.
3.2 Stage 1: Trajectory Extraction and Standardization
For a model $M$ and condition $A$, we construct probing sequences. At each time step $t$, we extract the hidden states of the last token across all layers at the decision-making cross-section immediately preceding the next assistant response generation:
@@ 3.2 Stage 1: Trajectory Extraction and Standardization
where $L$ is the number of layers and $d_{\ell}$ is the hidden dimension. As defined, these are aggregated into group-level representations $h_{t}^{(A,G)}\in\mathbb{R}^{d}$. Thus, for condition $A$ and group $G$, the sequence $\{h_{t}^{(A,G)}\}_{t=1}^{T_{A}}$ defines a trajectory in latent space.